OpenStudy (anonymous):

what is the square root of 1 over 9

OpenStudy (anonymous):

1 over 9.....

OpenStudy (radar):

\[\sqrt{1} \over \sqrt{9}\]

OpenStudy (radar):

Go ahead and take the square root of 1 Take the square root of 9.

OpenStudy (anonymous):

use this property :- \(\Huge {\sqrt\frac{{\color{Red} a}}{{\color{Blue} b}} } = \frac{\sqrt {\color{Red} a}}{\sqrt {\color{Blue} b}}\)